OpenStudy (anonymous):

19 + 2 x < 6 x + 20 Write answer in INTERVAL NOTATION.

OpenStudy (anonymous):

I got no clue yo

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Solve for x. Then you should get an inequality.

OpenStudy (anonymous):

i got x>−1/4 but tats wrong

OpenStudy (anonymous):

@theEric

OpenStudy (anonymous):

That's what I have. The inequality is saying that x can equal anything that is greater than -1/4. So to write it in interval notation you use ( for values that x cannot equal.

OpenStudy (anonymous):

how do i write it in interval notation x>−1/4 like this (-1/4+oo)

OpenStudy (anonymous):

\[19 + 2 x < 6 x + 20\\ 19<4x+20\\ -1<4x\\ -\frac{1}{4}<x\]

OpenStudy (anonymous):

\[(-\frac{1}{4},\infty)\]

OpenStudy (anonymous):

x is greater than -1/4. So that means infinity amounts of number greater than -1/4. You would write (-1/4,oo)

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

you guys are so helpful ;)

OpenStudy (theeric):

I don't think I saw anyone say exactly what I'm going to say, so I'll say it just in case! The paren "(" or ")" is used on that side of the interval if that boundary number isn't included in your interval. So, by saying \(\Large(\normalsize-\dfrac14, ...\) you say that \(-\dfrac14\) isn't part of your interval. That's correct, because \(x>-\dfrac14\), so \(x\) can't be \(-\dfrac14\). By the way, \(x\) just marks all the numbers that are accepted to be in the interval. I know you know that, but redundance can be good.

OpenStudy (theeric):

And you ALWAYS use the paren with infinity or negative infinity! Because, how can you include the number infinity? You really can't.. So, it's a paren.

OpenStudy (theeric):

Example: \(\left(-\infty,\infty\right)\)